Crew photos – April 10th

Sorry so late, this one was giving me trouble today.

Photo of the Day: HSO Ashok Narayanamoorthi on his first EVA on Mars


Sol Summary – April 10th

Crew 192 Sol 2 Summary Report

Sol 2

Summary Title: A Beautiful Day for EVAs

Author: Victoria Varone

Mission Status: It was a beautiful day for EVAs over here for Crew 192 on Mars. We continue to clean and organize things around the hab and have started getting more comfortable with EVAs, which will allow us to continue the search for fascinating natural wonders as well as our missing science experiments and crew member. The mission is going well.

Sol Activity Summary: The day started with a morning EVA, where three of our crew checked out the Marble Ritual site and surrounding area. They discovered some interesting geological structures and were particularly fascinated by a view of snow-capped mountains from a dusty, sandy red desert. After exploring, the crew returned to the hab, examined one of the hab’s water tanks for future cleaning, and headed inside for lunch.

After a great afternoon session of debriefing, brainstorming, and improvising for additional science experiments, the remaining crew members headed out for another EVA, coming back with excited descriptions of their hike in Kissing Camels Ridge. They also had some insight into the length of time of EVAs, noting that the amount of hiking they did in their 2.5 hour EVA expended just enough of their energy without causing serious dehydration. They’ve advise the rest of us to limit high-impact EVAs to about two and a half hours for safety.

We’re now resting and planning for tomorrow’s activities and may end the night with a space-based movie.

Look Ahead Plan: The goal is continue stretching our EVAs a little further as we keep searching for our missing science experiments and crew member. Based on tonight’s hiking EVA, we also plan to pay special attention to crew health on EVAs, making sure everyone is healthy and comfortable and no one becomes dehydrated and ill. A little attention is also being put towards possible activities for our Yuri’s Night celebration.

Anomalies in Work: None

Weather: The weather was beautiful again today, the skies started out clear and the weather cool, which worked well for the first EVA. It became overcast from the afternoon on, which allowed our second EVA crew members to explore the sites with different skies overhead.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: None

EVA Report – April 10th

EVA Report

EVA #2 – 10Apr2018 – Sol 2

Crew members: Richard Blakeman, Ashok Narayanamoorthi, Victoria Varone

Objectives:
– Further explore Marble Ritual site
– Document EVA
– Examine third water tank for state of cleanliness and usability.

Operations:
0815 Begin EVA
0830 Arrive at Marble Ritual
0900 Explore surrounding areas
1000 Return to hab site
1015 Examine third hab water tank
1030 Head back inside hab

Narrative:
Today was the crew’s second EVA of our mission. Three crew members (out of 5) went to explore a nearby site and surrounding areas, obtaining a lot of documentation of interesting geologic structures and formations and searching for possible signs of life in the reddish dirt. To head back to the hab, we traveled a little further with our vehicles to check out the landscape and utilized a turning point to turn back around. Upon arriving back at the hab, the rest of our crew inside requested that we take a quick survey of our third water tank to prepare it for being moved and cleaned. After that, we went back into the hab lunch.

Vehicles
1. Spirit Rover
2. Opportunity Rover
3. Blue ATV #3

______________

EVA Report

EVA #3 – 10Apr2018 – Sol 2

Crew members: Joe Dituri, Andreaa Radulescu

Objectives:
– Explore a new site
– Search for missing science experiments

Operations:
1630 – Entering the airlock
1635 – Move third water tank to RAM
1645 – Lock RAM and proceed to EVA
1730 – Arrive at Camel Ridge
1830 – Leave Camel Ridge
1900 – Arrive back at the hab

Narrative:
Today’s afternoon EVA was an opportunity to explore the Martian surface a little further than we did the day before. First, we completed a task for the crew, which was to move the third external water tank into our RAM for cleaning during our next few EVAs. After locking it inside the RAM, we headed out onto the Martian surface to an area called Kissing Camel Ridge, where we hiked for the majority of our EVA. We ended up finding a small badge of our missing crew member and decided to hike further in the area. The hike used up a lot of our energy, so we capped off our EVA at a little over two hours and headed back to the hab.

Vehicles
1. Red ATV
2. Blue ATV #3

Operations Report – April 10th

Crew 192 Operations Report 10Apr2018

SOL 2

Name of person filing report: Victoria Varone

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): 13 hours

Solar — SOC 100%

Diesel – 70%

Propane – 68% (telemetry reported) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – N/A

Water (trailer) – N/A
Water (static) – 1020 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133303.1

Toilet tank emptied: No. Last emptied April 9.

ATV’s Used: Blue #3, Red #1, and Blue #2.

Oil Added? No

ATV Fuel Used: 25%

# Hours the ATVs were Used today: 4

Notes on ATVs:

Deimos rover used: No
Hours: 0
Beginning charge: N/A
Ending charge: N/A
Currently charging: Yes

Spirit rover used: Yes
Hours: 2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 2
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 0
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

HabCar used and why, where? No

General notes and comments: A good day overall, the crew is doing very well.

Summary of internet: 569 MB remaining at 9:30PM.

Summary of suits and radios: Suits all charging, 4 one-piece suits used today. Radios working.

Summary of Hab operations: Changed out air filter, constructed minor and non-permanent communications modifications for radio clips in spacesuits, secured step outside of the side airlock.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None.

Summary of RAM operations: The third large external water tank on the trailer was rolled into the RAM for cleaning during additional EVAs.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support: The hot water faucet in the hab kitchen appears to be faulty, it was difficult to shut off after use. To avoid the loss of water, we’re not going to use it.

Journalist Report – April 10th

Journalist Report – 10 April 2018

Author: Richard Blakeman

We started this morning on a high note as I led crew members Ashok Narayanamoorthi and Victoria Varone out on our first Martian EVA. We specifically scheduled the day’s first EVA in the morning hours to avoid the strong afternoon sun, as crew members Joe Dituri and Andreaa Radulesco reported getting a little hot during their first EVA yesterday afternoon.

This allowed us to stay out a bit longer and explore further into the area, where we documented many different geologic formations, colors, and structures. We were very excited to come across erosion that was likely caused by liquid water flow, as well as some interesting animals tracks, both of which suggest possible life on Mars.

Our second EVA took place late in the afternoon, where crew members Joe Dituri and Andreaa Radulesco began by moving our third large water tank into the RAM for cleaning, which will be continued tomorrow. They locked up the RAM and headed out to explore under an overcast, cloudy sky.

Today was also our first day conducting two EVAs where we alternated crew members to allow for the best possible use of energy and time and to allow the whole crew enough time to explore the Red Planet.

Crew morale remains outstanding as we’ve been finding many ways to occupy our time, including minor fixes and adjustments around the hab, experimenting with cooking our dehydrated supply of food, brainstorming EVAs, and playing games. We also just found the hab’s stash of movies which will supply a wonderful additional dose of entertainment for our movie-loving crew.

Blue Crew 192 will continue their adventures on Mars tomorrow.

Richard Blakeman

Greenhab Report – April 10th

GreenHab Report
Name: Andreea Radulescu
Date: 10 Apr 2018

Environmental control:

Ambient with door open
Shade cloth on
Fan on (15:15~18:30)

Average temperatures: 30.6℃
Low 14.6℃
High 45.7℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 5 gal

Time(s) of watering for crops: Once in the morning around 9am

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: Picked few tomatoes and the crew ate them during brunch.

Support/supplies needed: None

Commander Report – April 10th

Mission Commander’s Report – SOL 2 – 10 April 2018

Author: Joe Dituri

With our morning EVA complete, we are formulating a plan for future EVAs and an effort to find our science suite and remnants of our science advisor. On our evening EVA, we were super excited to find a plaque card with our fallen science officer’s name on it. This renewed our optimism to find our science module in parts and salvage some of the science. The doctor continues to weigh the entire crew before and after EVA. So far most of the crew loses 1-3 lbs of body water per EVA depending upon duration. The doctor has continued to push fluids on the crew at all times for optimal health. He is also doing Skeletal muscle measurements for glycogen storage. We are performing isometric handgrip study prior to EVA to evaluate dexterity, numbness and comfort using EVA gloves. Yesterday we also started the well-being surveys. XO has started study to mitigate fogging of visors. Tomorrow the XO will submit an intermittent science report which will further detail science.

The crew played “Cards Against Humanity” last night. It served the purpose of lightening the mood after a hard day’s work. In the morning everyone woke refreshed and ready to start the day. The morning EVA served its purpose to indoctrination the three-person team for EVAs. The team explored the ridge to the east and found evidence of water as well as animal tracks.

The crew has worked up a plan to clean the water tank on the trailer. We took some of the EVA time to assess what was required to clean. We intend on rolling itinto the RAM tomorrow AM.

On the lighter side…some of the crew with longer hair compared dry shampoos and leave in conditioners and agreed to exchange product for a direct comparison taking into account a decreased ability to use water. Those participants who have insufficient hair to participate in the cross functional study will act as control group. Additionally hair constraint methods were discussed to reduce hair from being caught in the helmet

The crew took a great picture this morning with our red noses one and all to take a stand against childhood hunger. We were thinking about sending starving children some dehydrated food. Spirits are high and the crew is in a great mood.

LEARNED – We learned that the one-piece EVA suits have less mobility in the forward bending direction but they have different pressure points than the two-piece suit.

EVAs that involve a large amount of hiking should be limited to two or maybe two and a half hours unless there is supplemental water.

South of Kissing Camel Ridge, comms go out and you cannot get in touch with the hab.

IMPRESSED – I was impressed with the XO (Richard – Blue 01) as he led the EVA team and noticed the communication microphones were ill fitted. He devised a non – permanent alteration for the microphones which mimics a boom microphone and increased the effective communication. BRAVO ZULU to the XO.

POA&M – (Plan of Action and Milestones) – We are going to attempt to explore the Kissing Camel area where the science module plaque was found. We now have a training schedule for the crew to learn about decompression in space suits and psychological crew selection as well as radiation exposure and mitigation factors. Yuri’s night is the 12th of April and we made an attempt at planning some fun albeit limited.

~Dituri Sends (Blue 00)